A proposed cost-saving device has an installed cost of $810,000. The device will be used in a five-year project but is classified as three-year MACRS property for tax purposes. The required initial net working capital investment is $85,000, the marginal tax rate is 25 percent, and the project discount rate is 10 percent. The device has an estimated Year 5 salvage value of $130,000. What level of pretax cost savings do we require for this project to be profitable?
Answer #1
1] Initial cash flow = -810000-85000 = $         -8,95,000
2] PV of depreciation tax shields:
Year Depreciation % Depreciation Tax shield at 25% PVIF at10% PV at 10%
1 33.33 $           2,69,973 $                 67,493 0.90909 $         61,358
2 44.45 $           3,60,045 $                 90,011 0.82645 $         74,389
3 14.81 $           1,19,961 $                 29,990 0.75131 $         22,532
4 7.41 $              60,021 $                 15,005 0.68301 $         10,249
0.62092
Total 100 $           8,10,000 $            2,02,500 3.79079 $      1,68,528
PV of depreciation tax shields = $           1,68,528
3] PV of terminal non operating cash flows:
After tax salvage value = 130000*(1-25%) = $              97,500
Recovery of NWC $              85,000
Terminal non operating cash flow $           1,82,500
PV of terminal non operating cash flow = 182500*0.62092 = $           1,13,318
4] Sum PV of known cash flows = -895000+168528+113318= $        -6,13,154
For the project to be profitable the NPV should be
at the least .
For 0 NPV, the PV of the after tax cost savings should be at least 613154/3.79079 = $           1,61,748
Hence, pretax cost savings required = 161748/(1-25%) = $           2,15,665
[Answer]